BCF forecasts a 25% market share of total 1kg blocks of chocolate sold in Otago in December 2019 in Springfield. Reliable indFor this question ignore the scenario proposed in Question (b). Calculate the market size and market share variances for BCF for the month of December 2019.
Answer #1

Market size variance = (Actual industry sales in units - budgeted industry sales in units) * budget market share percentage * budgeted average unit contribution margin

Actual industry sales = 10,800 + 3,600 + 7,200 = 21,600

Budgeted industry sales = 7,200 + 4,800 + 4,000 = 16,000

budget market share percentage = 25%

budgeted average unit contribution = ($ 18,000 + $ 19,200 + $ 20,000)/(7,200+4,800+4,000)

= 3.575

Market size variance = (21,600 - 16,000)*25%*3.575

= 5,005

Market share variance = (actual market share percentage - budgeted market share percentage) * Actual industry sales in units * budgeted average unit contribution margin

= 12.5% * 21,600 * 3.575 = 9652.5